OpenStudy (dopeeaf):

Check my answer? Medal and fan! Which is the equation of a 20 cm diameter circle with its center at (3, -6)? (x+3)^2+(y-6)^2=100 (x-3)^2+(y+6)^2=400 (x-3)^2+(y+6)^2=100

OpenStudy (michele_laino):

here we have to apply the definition of a circumference, namely: "it is the set of all points whose distance from the center is equal to the radius". Now using the Theorem of Pitgora, we can write such statement as follows: \[{\left( {x - 3} \right)^2} + {\left( {y - \left( { - 6} \right)} \right)^2} = {20^2}\]

OpenStudy (michele_laino):

oops.. I have made a typo: \[{\left( {x - 3} \right)^2} + {\left( {y - \left( { - 6} \right)} \right)^2} = {\left( {\frac{{20}}{2}} \right)^2}\]

OpenStudy (michele_laino):

so, your answer is correct!